OpenStudy (ledah):

Simplify: 9x^3+6x^2+3x divided by 3x

OpenStudy (shane_b):

\[\frac {9x^3+6x^2+3x}{3x}=\frac {9x^3}{3x}+\frac{6x^2}{3x}+\frac{3x}{3x}\]Do you know how to divide exponents?

OpenStudy (ledah):

if your dividing exponents aren't you supposed to subtract them?

OpenStudy (shane_b):

Exactly. Divide the coefficients normally and subtract the exponents.

OpenStudy (ledah):

3x^2+2x+1 ?

OpenStudy (shane_b):

Perfect :)

OpenStudy (ledah):

awesome thanks

OpenStudy (shane_b):

yw, good luck!